# Neural Quantum State Study of Fracton Models

### v1

@@ Abstract
Fracton models host unconventional topological orders in three and higher dimensions and provide promising candidates for quantum memory platforms. Understanding their robustness against quantum fluctuations is an important task but also poses great challenges due to the lack of efficient numerical tools. In this work, we establish neural quantum states (NQS) as new tools to study phase transitions in these models. Exact and efficient parametrizations are derived for three prototypical fracton codes — the checkerboard and X-cube model, as well as Haah’s code — both in terms of a restricted Boltzmann machine (RBM) and a correlation-enhanced RBM. We then adapt the correlation-enhanced RBM architecture to a perturbed checkerboard model and reveal its strong first-order phase transition between the fracton phase and a trivial field-polarizing phase. To this end, we simulate this highly entangled system on lattices of up to 512 qubits with high accuracy, representing a cutting-edge application of variational neural-network methods. Our work demonstrates the remarkable potential of NQS in studying complicated three-dimensional problems and highlights physics-oriented constructions of NQS architectures.
@@ 1 Introduction
NQS represent a modern framework for addressing the ground state problem of strongly correlated systems [31] that is not inherently limited by dimensionality or the sign problem [32, 33]. The exponential growth of the Hilbert space is avoided by restricting the ground state search to a smaller subspace parameterized by neural networks. NQS can be considered a subclass of variational Monte Carlo (VMC) [34] techniques. The restricted Boltzmann machine (RBM) is the prototypical shallow NQS architecture [31, 35, 36, 37] and enjoys much popularity due to its relation with statistical physics, simple closed-form expressions, and connection to tensor network states [38, 39]. Although previous works have shown the RBMs’ ability to represent highly entangled systems and topological orders [40, 33, 41, 42, 43], they typically work in low dimensions or rely on the exactly solvable limit of topological stabilizer codes.
@@ 2.2 X-cube Model
For the X-cube model, a qubit is placed on every link of the three-dimensional cubic lattice. Then, we define $A_{\mathcal{C}}\coloneqq\prod_{i\in\mathcal{C}}\sigma_{i}^{x}$ for every cube $\mathcal{C}$, where the tensor product involves all $12$ qubits on the edges of the cube. Further, to each vertex $\mathcal{V}$ we assign three different crosses: $\mathcal{S}^{x}=\set{n,j,p,m}$, $\mathcal{S}^{y}=\set{n,i,p,k}$, and $\mathcal{S}^{z}=\set{k,j,i,m}$, the indices of which are marked in Figure 1(b). In this notation, $S^{\mu}$ is the star orthogonal to direction $\mu\in\set{x,y,z}$. Then, for every vertex $\mathcal{V}$ we define the star operators as $B_{\mathcal{S}^{\mu}}\coloneqq\prod_{i\in\mathcal{S}^{\mu}}\sigma_{i}^{z}$. Finally, the Hamiltonian of the X-cube model can be written as
@@ 3 Neural Quantum States
This is done by estimating the gradient of $E(\boldsymbol{\theta})$ with Monte Carlo samples from the Born distribution $|\psi_{\boldsymbol{\theta}}|^{2}$ and updating the parameters into the direction of steepest descent using stochastic reconfiguration, first introduced by [54]. Repeating this parameter update iteratively then leads to a local, ideally global, minimum of the energy landscape $E(\boldsymbol{\theta})$. Hence, neural quantum states (NQS) can be considered a subclass of variational techniques. Appendix B contains a more detailed description of the NQS optimization.
@@ 4.2 X-cube Model
Similar to the checkerboard model, one hidden unit is connected locally to every $\mathcal{S}^{x},\mathcal{S}^{y}$ and $\mathcal{S}^{z}$ cross corresponding to the $\sigma^{z}$-type stabilizer generators, as shown in Figure 8. We directly set all visible biases to zero to arrive at the following RBM ansatz:
@@
Figure 8: The hidden units (right) are connected locally to the three different crosses $\mathcal{S}^{x},\mathcal{S}^{y}$ and $\mathcal{S}^{z}$ at each corner in the X-cube model (left). The three qubits in $\mathcal{C}$ associated with the corner in the bottom left are colored red.
@@
To find the expression of the X-cube ground state in terms of a cRBM, we introduce cross correlators for all $\mathcal{S}^{\mu}$ and connect one hidden unit to each of them. After setting all visible biases to zero, we get
@@
Again, the cRBM parametrization contains less non-zero parameters, i.e. $6N$, compared to the RBM ansatz with $12N$ parameters. At this point, the general scheme for constructing an exact and efficient cRBM parametrization for any CSS code should become clear: Introduce a correlator feature for every $\sigma^{z}$-type stabilizer generator $T^{z}_{i}$ and locally connect a hidden unit to each of them with corresponding weights $b_{i}=\mathrm{i}\frac{\pi}{4}$, $W_{i}=-\mathrm{i}\frac{\pi}{4}$ and visible biases set to zero. For a stabilizer code, all $\sigma^{x}$-type stabilizer generators $T^{x}_{i}$ commute with any $T^{z}_{j}$, so they must share an even number of qubits. Hence, the $T^{x}_{i}$ generator conditions are directly satisfied.
@@ 4.3 Haah’s Code
The linear scaling in system size of all derived exact parametrizations can be made constant by further imposing translational symmetries. For instance, two hidden units for every cube are capable of representing the unperturbed checkerboard model as indicated in Figure 6 and explicitly shown in Appendix A. Such knowledge of exact representations can guide the architecture design for perturbed fracton models where exact solutions become unavailable.
@@ 5 Simulation Overview
Moreover, we introduce a transfer learning protocol that carries over samples and parameters of a trained NQS to the optimization for the next physical parameter, namely, a nearby value of the magnetic field $\vec{h}$. By comparing the results for transfer learning along increasing and decreasing fields, the method detects hysteresis effects and provides insights into the nature of the underlying phase transitions.
Sampling schemes like the Metropolis-Hasting algorithm can suffer from many issues such as long auto-correlation times, for instance. We make use of a custom update rule including, for example, cube flips (see Appendix B), perform sampling on $O(1000)$ parallel chains convergence, and carefully analyze the convergence of our simulations by monitoring behaviors of the acceptance ratio of our update rule, the split-$\hat{R}$ value [57], and the V-score [58] in Appendix D.
@@ 6 Benchmarking on Small Lattices
We benchmark the performance of the cRBM on a checkerboard model of size $4\times 2\times 2$ with PBCs subject to uniform magnetic fields along different directions. Such a system size remains exactly diagonalizable, hence we can compare with the exact results.
First, we compare the performance of several common NQS architectures on the pure checkerboard model in the absence of magnetic fields. Figure 10 shows the training progress for an FFNN with two hidden layers containing 8 hidden units each, a regular RBM with 4 hidden units, a translation-invariant RBM with $\alpha=1/4$, a symmetric FFNN that first constructs 4 translation-invariant features (as in a symmetric RBM) which are then processed in another FFNN-layer before being accumulated, and a translation-invariant cRBM architecture with $\alpha=1/4$ containing only additional cube correlators. Clearly, the cRBM is the only architecture that converges to the ground state, while all other models remain in some initial local minimum at approximately $E_{0}/2$. The initial parameters are sampled independently from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ with zero mean and small standard deviation $\sigma\sim 10^{-2}$, a common practice machine learning. This implies that the initial state described by any network with a $\cosh$-activation function approximately corresponds to the equal-weighted superposition of all configurations |$\ket{\psi_{\boldsymbol{\theta}\approx 0}}_{z}\approx 2^{-\frac{N}{2}}\otimes_{i}(\ket{+z}_{i}+\ket{-z}_{i})=\otimes_{i}\ket{+x}_{i}$. Consequently, all $A_{\mathcal{C}}$ stabilizer conditions are approximately satisfied, resulting in an initial energy of $E(\theta_{0})\approx E_{0}/2$. Apparently, the transfer from an equal weighted superposition of all states to a "fracton soup", the equal weighted superposition of states which also satisfy the $B_{\mathcal{C}}$ stabilizer conditions, comes with a temporary increase of the variational energy. This causes all models except the cRBM to remain in the initial local minimum for a significant number of training iterations. The cube correlators included in the cRBM transform the loss landscape in such a way that the initial local minimum is flattened or even removed. This highlights the significant advantage achieved through incorporating domain knowledge by including custom correlator features.
@@
Figure 10: Comparison of different model architectures for learning the ground state of the pure Checkerboard model on a $4\times 2\times 2$ lattice. The parameter counts for the different models are 208 for the FFNN, 84 for the RBM, 35 for the symmetric RBM, 88 for the symmetric NN and 52 for the cRBM architecture. The ground state energy is indicated by the dashed grey line.
Next, relative energy errors of the trained cRBM for the checkerboard model Eq. 9 perturbed with different fields are shown in Figure 11. Implementation details and hyperparameter choices are contained in Appendix C. We observe a significant difference in accuracy between magnetic fields applied along the $x$- and $z$-directions, although the only difference between these two field configurations is a rotation from the $x$- to the $z$-basis. Hence, this discrepancy must be due to the choice of basis made when expanding the wave function, the only step where the $x$- and $z$-directions are treated differently. Of course, the choice of basis inherently determines the form of the loss landscape $E(\boldsymbol{\theta})$, and hence the difficulty of the optimization problem. In particular, however, it determines the initial state in the following sense: As mentioned before, the initial parameters are typically sampled independently from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ with zero mean and small standard deviation $\sigma\sim 10^{-2}$. In the $z$-basis, this implies that the initial state described by an RBM or cRBM approximately corresponds to the equal-weighted superposition of all configurations |$\ket{\psi_{\boldsymbol{\theta}\approx 0}}_{z}\approx 2^{-\frac{N}{2}}\otimes_{i}(\ket{+z}_{i}+\ket{-z}_{i})=\otimes_{i}\ket{+x}_{i}$. On the other hand, when working in the $x$-basis this parameter initialization results in an initial state that is close to the $z$-polarized state |$\ket{\psi_{\boldsymbol{\theta}\approx 0}}_{x}\approx 2^{-\frac{N}{2}}\otimes_{i}(\ket{+x}_{i}+\ket{-x}_{i})=\otimes_{i}\ket{+z}_{i}$. Together with the results in Figure 11, this suggests that the performance for some applied magnetic field might be improved by initializing the NQS close to the corresponding polarized state.
@@
With that motivation in mind, we introduce a simple transfer-learning protocol: First, the NQS is optimized in the presence of strong magnetic fields, which can be done for all field directions with high accuracy (see Figure 11). Then, the NQS is trained for the next highest field strength with the previously optimized NQS as the initial state. We achieve this by taking the learned parameters and, whenever full summation over the Hilbert space is not possible due to larger system sizes, also the final states of the Markov chains as the starting point for the NQS optimization with the new magnetic field. This process is repeated until some smallest field strength, usually zero, is reached. Hence, we not only initialize the NQS close to the polarized state but we also recycle learned features by transferring the network parameters. See Appendix B for additional details.
@@ 7 Field-driven Phase Transitions
We now move on to larger systems of size $L\times L\times L$, $L\in\set{4,6,8}$, corresponding to 64, 216 and 512 qubits, respectively. The low parameter count of the cRBM, i.e. 5195 for $L=8$, makes these computations feasible even with limited computational resources. The observables and gradients are estimated with samples obtained from the Metropolis algorithm; in addition to single spin flips, our update rule includes cube and non-contractible loop flips, see Appendix B and C for further technical details.
@@
Moreover, although the benchmark results in Fig. 12 suggest better performance when sweeping from strong to weak fields (right-left), we also include sweeps from weak to strong fields (left-right) for the following reason: Left-right sweeps, in particular along the $x$-direction, still result in a reasonably high energy accuracy and comparing the two sweep directions will help uncover a hysteresis.
@@
As shown in Figure 14 for lattice sizes $L=4,6,8$, the energy and the magnetization hystereses become more profound as revealed by comparing the left-right and right-left transfer learning results. When sweeping from right to left, the NQS settles in a meta-stable phase resulting in higher energy for fields slightly weaker than the critical value $h_{\mathrm{crit}}$. When the field strength is decreased further, the state becomes unstable and collapses into a new state with lower energy. This sudden jump coincides with a discontinuous decrease of the magnetization. For left-right sweeps we observe similar behaviour, but for system sizes $L>4$ the NQS remains in the meta-stable phase for all observed field strengths. This confirms that the perturbed checkerboard model indeed experiences a strong first-order phase transition. We estimate a critical field $h_{\mathrm{crit}}\approx 0.445$ from the intersection points of energy curves and the middle of the magnetization hysteresis.

### v2

@@ Abstract
Fracton models host unconventional topological orders in three and higher dimensions and provide promising candidates for quantum memory platforms. Understanding their robustness against quantum fluctuations is an important task but also poses great challenges due to the lack of efficient numerical tools. In this work, we establish neural quantum states (NQS) as new tools to study phase transitions in these models. Exact and efficient parametrizations are derived for three prototypical fracton codes — the checkerboard and X-cube model, as well as Haah’s code — both in terms of a restricted Boltzmann machine (RBM) and a correlation-enhanced RBM. We then adapt the correlation-enhanced RBM architecture to a perturbed checkerboard model and reveal its strong first-order phase transition between the fracton phase and a trivial field-polarizing phase. To this end, we simulate this highly entangled system on lattices of up to 512 qubits with high accuracy, representing a cutting-edge application of variational neural-network methods. In addition, we reproduce the phase transition of the X-cube model previously obtained with quantum Monte Carlo and high-order series expansion methods. Our work demonstrates the remarkable potential of NQS in studying complicated three-dimensional problems and highlights physics-oriented constructions of NQS architectures.
@@ 1 Introduction
NQS represent a modern technique for addressing the ground state problem of strongly correlated systems [31] that is not inherently limited by dimensionality or the sign problem [32, 33]: The quantum wave function is parameterized in terms of a neural network, thereby avoiding the exponential growth of the Hilbert space by restricting the ground state search to a smaller parameterized subspace. Observables and gradients can then be estimated using Monte-Carlo sampling, which places NQS in the variational Monte Carlo (VMC) [34] framework.
The restricted Boltzmann machine (RBM) is the prototypical shallow NQS architecture [31, 35, 36, 37] and enjoys much popularity due to its relation with statistical physics, simple closed-form expressions, and connection to tensor network states [38, 39]. Although previous works have shown the RBMs’ ability to represent highly entangled systems and topological orders [40, 33, 41, 42, 43], they typically work in low dimensions or rely on the exactly solvable limit of topological stabilizer codes.
@@ 2.2 X-cube Model
For the X-cube model, a qubit is placed on every link of the three-dimensional cubic lattice. Then, we define $A_{\mathcal{C}}\coloneqq\prod_{i\in\mathcal{C}}\sigma_{i}^{x}$ for every cube $\mathcal{C}$, where the tensor product involves all $12$ qubits on the edges of the cube. Further, to each vertex $\mathcal{V}$ we assign three different stars: $\mathcal{S}^{x}=\set{n,j,p,m}$, $\mathcal{S}^{y}=\set{n,i,p,k}$, and $\mathcal{S}^{z}=\set{k,j,i,m}$, the indices of which are marked in Figure 1(b). In this notation, $S^{\mu}$ is the star orthogonal to direction $\mu\in\set{x,y,z}$. Then, for every vertex $\mathcal{V}$ we define the star operators as $B_{\mathcal{S}^{\mu}}\coloneqq\prod_{i\in\mathcal{S}^{\mu}}\sigma_{i}^{z}$. Finally, the Hamiltonian of the X-cube model can be written as
@@ 3 Neural Quantum States
This is done by estimating the gradient of $E(\boldsymbol{\theta})$ with Monte Carlo samples from the Born distribution $|\psi_{\boldsymbol{\theta}}|^{2}$ and updating the parameters into the direction of steepest descent using stochastic reconfiguration, first introduced by [54]. Repeating this parameter update iteratively then leads to a local, ideally global, minimum of the energy landscape $E(\boldsymbol{\theta})$. Hence, neural quantum states (NQS) are a subclass of variational quantum states characterized by parameterizing the wave function in terms of a neural network. Appendix B contains a more detailed description of the NQS optimization.
@@ 4.2 X-cube Model
Similar to the checkerboard model, one hidden unit is connected locally to every $\mathcal{S}^{x},\mathcal{S}^{y}$ and $\mathcal{S}^{z}$ star corresponding to the $\sigma^{z}$-type stabilizer generators, as shown in Figure 8. We directly set all visible biases to zero to arrive at the following RBM ansatz:
@@
Figure 8: The hidden units (right) are connected locally to the three different stars $\mathcal{S}^{x},\mathcal{S}^{y}$ and $\mathcal{S}^{z}$ at each corner in the X-cube model (left). The three qubits in $\mathcal{C}$ associated with the corner in the bottom left are colored red.
@@
To find the expression of the X-cube ground state in terms of a cRBM, we introduce star correlators for all $\mathcal{S}^{\mu}$ and connect one hidden unit to each of them. After setting all visible biases to zero, we get
@@
Again, the cRBM parametrization contains less non-zero parameters, i.e. $6N$, compared to the RBM ansatz with $12N$ parameters. For completeness, we point to [55] for an alternative parametrization of the X-cube ground state. At this point, the general scheme for constructing an exact and efficient cRBM parametrization for any CSS code should become clear: Introduce a correlator feature for every $\sigma^{z}$-type stabilizer generator $T^{z}_{i}$ and locally connect a hidden unit to each of them with corresponding weights $b_{i}=\mathrm{i}\frac{\pi}{4}$, $W_{i}=-\mathrm{i}\frac{\pi}{4}$ and visible biases set to zero. For a stabilizer code, all $\sigma^{x}$-type stabilizer generators $T^{x}_{i}$ commute with any $T^{z}_{j}$, so they must share an even number of qubits. Hence, the $T^{x}_{i}$ generator conditions are directly satisfied.
@@ 4.3 Haah’s Code
The linear scaling in system size of all derived exact parametrizations can be made constant by further imposing translational symmetries. For instance, a translation invariant RBM architecture with two hidden units is capable of representing the ground state of the unperturbed checkerboard model exactly, as shown in Figure 6 and Appendix A. Such knowledge of exact representations can guide the architecture design for perturbed fracton models where exact solutions become unavailable.
@@ 5 Simulation Overview
Moreover, we introduce a transfer learning protocol that carries over samples and parameters of a trained NQS to the optimization for the next physical parameter, namely, a nearby value of the magnetic field $\vec{h}$, see Appendix B. By comparing the results for transfer learning along increasing and decreasing fields, the method detects hysteresis effects in the checkerboard model and provides insights into the nature of the underlying phase transitions. We verified this method on the X-cube model, for which existing quantum Monte Carlo and high-order series expansion results predict a strong first-order transition, see Appendix E.2.
Sampling schemes like the Metropolis-Hasting algorithm can suffer from many issues such as long auto-correlation times, for instance. We make use of a custom update rule including, for example, cube flips (see Appendix B), perform sampling on $O(1000)$ chains in parallel, and carefully analyze the convergence of our simulations by monitoring behaviors of the acceptance ratio of our update rule, the split-$\hat{R}$ value [58], and the V-score [59] in Appendix D.
@@ 6 Benchmarking on Small Lattices
We benchmark the performance of the cRBM on a checkerboard model of size $4\times 2\times 2$ with PBCs subject to uniform magnetic fields along different directions. Such a system size remains exactly diagonalizable, hence no sampling is required and we can compare with the exact results.
First, in Figure 10 we compare the performance of several common NQS architectures on the pure checkerboard model in the absence of magnetic fields and for different combinations of hyperparameters, namely the standard deviation of the parameter initialization and the learning rate. We make use of the following architectures: an FFNN with two hidden layers containing 8 hidden units each, a regular RBM with 16 and 8 hidden units, a translation-invariant RBM with 8 and 4 hidden units, a symmetric FFNN that first constructs 4 translation-invariant features (as in a symmetric RBM) which are then processed by another layer consisting of 4 hidden units before being accumulated, a Jastrow ansatz, and a translation-invariant cRBM with $\alpha=1/4$ containing only additional cube correlators. The Jastrow wave function is modeled as $\psi_{\mathrm{Jastrow}}(\boldsymbol{\sigma})=\exp(\sum_{i\neq j}\sigma_{i}W_{ij}\sigma_{j})$ with a symmetric matrix $W$ that is treated as lower triangular. All types of RBM architectures except RBM(8) contain enough hidden units to learn the exact representations derived in Section 4.
((a)) $\sigma=0.01,\ \mathrm{lr}=0.01$
((b)) $\sigma=0.01,\ \mathrm{lr}=0.1$
((c)) $\sigma=0.1,\ \mathrm{lr}=0.01$
((d)) $\sigma=0.1,\ \mathrm{lr}=0.1$
@@
Figure 10: Comparison of different model architectures for learning the ground state of the pure Checkerboard model on a $4\times 2\times 2$ lattice. The complex parameters are sampled independently from a centered complex Gaussian with given standard deviation. The number in parentheses corresponds to the number of hidden units. The parameter counts for the different models are 208 for the FFNN, 288/152 for the RBM(16)/(8), 137/69 for the symmetric RBM(8)/(4), 88 for the symmetric NN, 120 for the Jastrow and 52 for the cRBM architecture. The ground state energy is indicated by the dashed grey line.
The initial complex-valued parameters are sampled independently from a complex Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ with zero mean and standard deviation $\sigma$, a common practice in machine learning. For sufficiently small $\sigma$, this implies that the initial state described by any network with a final $\cosh$-activation function approximately corresponds to the equal-weighted superposition of all configurations |$\ket{\psi_{\boldsymbol{\theta}\approx 0}}_{z}\approx 2^{-\frac{N}{2}}\otimes_{i}(\ket{+z}_{i}+\ket{-z}_{i})=\otimes_{i}\ket{+x}_{i}$. Consequently, all $A_{\mathcal{C}}$ stabilizer conditions are approximately satisfied, resulting in an initial energy of $E(\theta_{0})\approx E_{0}/2$. This state, however, appears to be a strong local minimum for all but the cRBM architecture, as can be seen in Figure 10(a). The challenge of transitioning from an equal weighted superposition of all states to a fracton soup, the equal weighted superposition of states which also satisfy the $B_{\mathcal{C}}$ stabilizer conditions, is lifted for the cRBM by simplifying its loss landscape through the inclusion of the cube correlators.
By increasing the variance of the parameter initialization as in Figure 10(c), it is possible to move away from said local minimum, thereby allowing the (symmetric) RBM architectures with $\alpha=1$ to reach the ground state within the given number of training iterations. A larger variation in the initial parameters can result in unstable training and may lead to local minima later in the optimization, especially for larger learning rates as in Figure 10(d). Also, we refer to Figure 21 in Appendix E.1 which displays the training curves for uniform parameter initialization. Figure 10(c) additionally shows training curves for the (symmetric) RBM architectures with only half as many hidden units. Clearly, the reduced $\alpha$ value results in non-convergence of these RBM architectures within the given number of training steps, although both models still contain more parameters than the cRBM. In conclusion, while other network architectures are able to approximate the ground state successfully in some scenarios, the cRBM architecture displays significant advantage over the others: it requires the smallest number of training iterations to learn the ground state, it exhibits strong robustness to different hyperparameter choices, it typically reaches the lowest energy variance by the end of training, and it is the most parameter efficient parametrization of the considered candidates. This highlights the significant advantage achieved by incorporating domain knowledge.
@@
Next, relative energy errors of the trained cRBM for the checkerboard model perturbed with different fields (Eq. 9) are shown in Figure 11. Implementation details and hyperparameter choices are contained in Appendix C. We observe a significant difference in accuracy between magnetic fields applied along the $x$- and $z$-directions, although the only difference between these two field configurations is a rotation from the $x$- to the $z$-basis. Hence, this discrepancy must be due to the choice of basis made when expanding the wave function, the only step where the $x$- and $z$-directions are treated differently. Of course, the choice of basis inherently determines the form of the loss landscape $E(\boldsymbol{\theta})$, and hence the difficulty of the optimization problem. In particular, however, it determines the initial state in the following sense: Here, the initial parameters were sampled independently from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ with zero mean and small standard deviation $\sigma\sim 10^{-2}$. In the $z$-basis, this implies that the initial state described by an RBM or cRBM approximately corresponds to the equal-weighted superposition of all configurations |$\ket{\psi_{\boldsymbol{\theta}\approx 0}}_{z}\approx 2^{-\frac{N}{2}}\otimes_{i}(\ket{+z}_{i}+\ket{-z}_{i})=\otimes_{i}\ket{+x}_{i}$. On the other hand, when working in the $x$-basis this parameter initialization results in an initial state that is close to the $z$-polarized state |$\ket{\psi_{\boldsymbol{\theta}\approx 0}}_{x}\approx 2^{-\frac{N}{2}}\otimes_{i}(\ket{+x}_{i}+\ket{-x}_{i})=\otimes_{i}\ket{+z}_{i}$. Together with the results in Figure 11, this suggests that the performance for some applied magnetic field might be improved by initializing the NQS close to the corresponding polarized state.
With that motivation in mind, we introduce a simple transfer-learning protocol: First, the NQS is optimized in the presence of strong magnetic fields, which can be done for all field directions with high accuracy (see Figure 11). Then, the NQS is trained for the next highest field strength with the previously optimized NQS as the initial state. We achieve this by taking the learned parameters and, whenever full summation over the Hilbert space is not possible due to larger system sizes, also the final states of the Markov chains as the starting point for the NQS optimization with the new magnetic field. This process is repeated until some smallest field strength, usually zero, is reached. Hence, we not only initialize the NQS close to the polarized state but we also recycle learned features by transferring the network parameters. We would like to point out that this optimization scheme can also be understood as an implementation of variational neural annealing (VNA) [60]. A similar protocol has also been utilized in [61]. See Appendix B for additional details.
@@ 7 Field-driven Phase Transitions
We now move on to larger systems of size $L\times L\times L$, $L\in\set{6,8}$, corresponding to 6216 and 512 qubits, respectively. The low parameter count of the cRBM, i.e. 5195 for $L=8$, makes these computations feasible even with limited computational resources. The observables and gradients are estimated with samples obtained from the Metropolis algorithm; in addition to single spin flips, our update rule includes cube and non-contractible loop flips, see Appendix B and C for further technical details.
@@
Moreover, although the benchmark results in Fig. 12 suggest better performance when sweeping from strong to weak fields (right-left), we also include sweeps from weak to strong fields (left-right) for the following reason: Left-right sweeps, in particular along the $x$-direction, still result in a reasonably high energy accuracy, and comparing the two sweep directions will help uncover a hysteresis.
Refer to captionRefer to caption
@@
As shown in Figure 14 for lattice sizes $L=4,6,8$, the energy and magnetization hystereses become more profound, as revealed by comparing the left-right and right-left transfer learning results. When sweeping from right to left, the NQS settles in a meta-stable phase resulting in higher energy for fields slightly weaker than the critical value $h_{\mathrm{crit}}$. When the field strength is decreased further, the state becomes unstable and collapses into a new state with lower energy. This sudden jump coincides with a discontinuous decrease of the magnetization. For left-right sweeps we observe similar behavior, but for system sizes $L>4$ the NQS remains in the meta-stable phase for all observed field strengths. This confirms that the perturbed checkerboard model indeed experiences a strong first-order phase transition. We estimate a critical field $h_{\mathrm{crit}}\approx 0.44(1)$ from the intersection points of energy curves.